The price of shares of Bank of Florida at the end of trading each day for the last year followed the normal distribution. Assume there were 240 trading days in the year. The mean price was $42.00 per share and the standard deviation was $2.25 per share. (Round your answers to 2 decimal places.)

a-1. What is the probability that the end-of-day trading price is over $45.00? ______%
Answer #1

Solution :

    we are given that : The price of shares of Bank of Florida at the end of trading each day for the last year followed the normal distribution.

   Assume there were 240 trading days in the year. The mean price was $42.00 per share and the standard deviation was $2.25 per share.

Thus   Mean = mu =42.00    and Standard Deviation = σ 2.25

We have to find :

Part A. What is the probability that the end of day trading price is over $45.00?

That is we have to find : P( X > 45.00) = ..............?

Thus find z score for x = 45.00

z = rac{x-mu }{sigma }

5.00 -42.00

3.00 2.25

~ 1.33

Thus we get :

P( X > 45.00 ) = P( Z > 1.33)

P( X > 45.00 ) = 1 - P( Z < 1.33)

Look in z table for z = 1.3 and 0.03 and find corresponding area.

01 02 03 05 07 0.0 .5000 .5040 5080 .520 5160 5199 5239 5279 .5319 5359 0.1.5398 5438 5478 0.2 .5793 .5832 .5871 0.3 .6179 .6

Thus we get ; P( Z < 1.33 ) = 0.9082

Thus P( X > 45.00 ) = 1 - P( Z < 1.33)

        P( X > 45.00 ) = 1 - 0.9082

       P( X > 45.00 ) = 0.0918